if sarah can pain a fence in 4 hours and tim can paint the same fence in 8 hours how much would it take them to paint the fence toguether
Answer by nerdybill(7384)

if sarah can pain a fence in 4 hours and tim can paint the same fence in 8 hours how much would it take them to paint the fence toguether
.
Let x = time it takes for both to paint fence together
then
x(1/4 + 1/8) = 1
multiplying both sides by 8:
x(2 + 1) = 8
x(3) = 8
x = 8/3 hours
or
x = 2 and 2/3 hours
or
x = 2 hours and 40 minutes